# The impact of noise on the simulation of NMR spectroscopy on NISQ devices

### v1

@@ V.3 Effective decoherence rate
Figure 7: (a) NMR spectrum of 1,2,4-trichlorobenzene ($C_{6}H_{3}Cl_{3}$) obtained via exact diagonalization. The shaded area depicts the different spin couplings in ppm units, see Eq. 4. (b) Comparison of the NMR spectrum of the reduced Hamiltonian of 1,2,4-trichlorobenzene ($C_{6}H_{3}Cl_{3}$) simulated on IBM Perth (solid orange curve) with on-gate depolarizing noise model emulation (dashed blue curve). The shaded area depicts the effective decoherence rate Eq. (18) in ppm units.
Such an effective decoherence rate allows us to quantify the precision with which we can simulate the NMR spectrum on NISQ devices. In particular, in the case of 1,2,4-trichlorobenzene, see Figs. 4 and 5, the effective decoherence rate is much larger than the spin-spin coupling between spin 2 and 3 $\Gamma_{\rm eff}\gg J_{23}$, see Fig. 7(a). Consequently, we cannot resolve the energy splitting caused by $J_{23}$ interactions. We can therefore (in principle) neglect the $J_{23}$ terms in the Hamiltonian Eq. (4). Neglecting the coupling $J_{23}$ simplifies 1,2,4-trichlorobenzene to a nearest neighbor Hamiltonian. In the first-order Trotter expansion, each Trotter circuit then contains 6 CNOT gates (instead of 15 CNOTs). As a result, the simulated NMR spectrum of the reduced Hamiltonian is significantly improved, as can be seen by comparing Fig. 5 with Fig. 7. For completeness, we have also presented the simulation with depolarising noise, and we see that the resulting spectrum agrees well with the simulated one on IBMQ cloud.
@@ Acknowledgements.
This work was supported by the German Federal Ministry of Education and Research, through projects QSolid (13N16155) and Q-Exa (13N16065). We also want to thank Peter Pinski, Sonia Alvarez, and Julius Kleine Büning for insightful input and fruitful discussions.
@@ Appendix A Time evolution of Magnetization
In this appendix we look at the time evolution of the correlation function simulated on IBM quantum computing platforms. Fig. 8 depicts the correlation function $\langle S^{z}_{\rm tot}(t)S^{z}_{\rm tot}\rangle$ as a function of time, indicating the decay of correlation function with time. In the NMR experiment the spectrum is measured at sufficiently long times after the application of the magnetic field such that the system reaches a state in which average magnetization vanishes and hence the correlation function. As is shown in Fig. 8 there remain some small oscillations around zero at long times due to the limitation of number of Trotter steps. As the spectrum is obtained via the Fourier transform of the correlation function, see Eq. (12), the frequency interval is inversely proportional to the number of Trotter steps. Thereby we added additional zeros to the correlation functions in order to be able to obtain the spectrum on a finer frequency grid. Fig. 9 presents the comparison of the spectrum with and without adding such additional zeros for one of the simulations on IBM’s devices.
Figure 8: Time evolution of the correlation function of the cis-3-chloroacrylic acid, simulated on two different devices from IBM quantum computers, namely IBM Perth (orange curve) and IBM Nairobi (green curve), with 81 Trotter steps of size $\Delta t=0.01$.
@@ Appendix B The symmetrized NMR spectra
In this appendix, we discuss the symmetry breaking that occurs in the simulations as an artifact of noise of NISQ devices. In particular in Fig. 10, we present the magnetization in $y$ direction simulated on IonQ’s Aria. Note that due to the symmetry of the Hamiltonian of cis-3-chloroacrylic acid, see Eq. 2, we expect $\langle S^{0}_{y}\rangle=-\langle S^{1}_{y}\rangle$, such that the total magnetization in the $y$ direction vanishes. However, due to the presence of noise, such a symmetry is numerically broken. This is a manifestation of coherent errors [45] resulting in additional disordered effective Hamiltonian terms that break the mentioned symmetry. In particular, in additional numerical simulations we find that the spectroscopy on quantum computer is especially fragile against Hamiltonian disorder terms $\propto YZ$.
Figure 10: Time evolution of the magnetization in the $y$ direction for cis-3-chloroacrylic acid with Hamiltonian Eq. 2, simulated on IonQ Aria, with 61 Trotter steps of size $\Delta t=0.01$. The contribution from each spin to the magnetization (orange and blue curve) is compared with the total one (black curve).
Figure 11: Comparison of the symmetrized spectrum of the cis-3-chloroacrylic acid, simulated on IonQ Aria (dash-dotted blue curve) with 61 Trotter steps of size $\Delta t=0.01$, with exact diagonalization results (black solid curve).
@@ Appendix C The trotter error
For a Trotterized time evolution, it is advantageous to increase the trotter step size to be able to simulate longer times, and also to decrease the impact of noise, as discussed in Sec. V.3. However, by increasing the Trotter step size, we get an error from the Trotterization of the time evolution. Fig. 12 depicts the occurrence of such an error in the form of shifts of the resonances. Note that in the simulation with an on-gate depolarizing noise, we can reproduce such shifts to the actual resonances.
Figure 12: Simulation of NMR spectrum of cis-3-chloroacrylic acid on IBM Perth with 81 Trotter steps of size $\Delta t=0.02$ shown by orange curve, compared with exact results (black curve) as well as simulation with on-gate depolarizig noise model (dashed blue curve).
@@ Appendix D Additional results
In this appendix, we present further simulations of the NMR spectrum of cis-3-chloroacrylic acid on IBM’s processors. It is interesting to note that the deviations of the predicted resonances from the exact result varies in simulations performed on different days, which can be a manifestation of quasistatic coherent noise.
@@
Figure 13: Comparison of the simulation of NMR spectrum of cis-3-chloroacrylic acid on IBM Perth on three different days (shown with different colors) with exact results (black curve). For the simulation we performed a time evolution with 81 Trotter steps of size $\Delta t=0.01$.

### v3

@@ V Noise analysis
Figure 6: (a) NMR spectrum of 1,2,4-trichlorobenzene ($C_{6}H_{3}Cl_{3}$) obtained via exact diagonalization. The shaded area depicts the different spin couplings in ppm units, see Eq. 4. (b) Comparison of the NMR spectrum of the reduced Hamiltonian of 1,2,4-trichlorobenzene ($C_{6}H_{3}Cl_{3}$) simulated on IBM Perth (solid orange curve) with on-gate depolarizing noise model emulation (dashed blue curve). The shaded area depicts the effective decoherence rate Eq. 13 in ppm units.
In Fig. 6, we show the linewdith given by $\Gamma_{\rm eff}$ as compared to an actual run on a quantum computer for a three spin example, see Eq. 4. We see that there is again a good agreement. This example is quite important and interesting to understand how we can directly use this simple rate to make meaningful predictions for the results of our calculation and even use it to improve our calculations. As is clear from the fact that $\Gamma_{\rm eff}$ is an effective line width, it represents the limit of energy scales that can be resolved. In fig. 6. (a), we show the coupling $J_{23}$ and in fig. 6. (b) we see that the size of $\Gamma_{\rm eff}$ is orders of magnitudes larger than the spin coupling, $\Gamma_{\rm eff}>J_{23}$. Note that all this is known already before actually simulating the spectra. The coupling known from the input Hamiltonian and the rate $\Gamma_{\rm eff}$ can be calculated using the parameters provided by the quantum computing providers. Therefore we can directly simplify the problem by neglecting $J_{23}$ which reduces the gates depth of the circuit. Neglecting the coupling $J_{23}$ simplifies 1,2,4-trichlorobenzene to a nearest neighbor Hamiltonian. In the first-order Trotter expansion, each Trotter circuit then contains 6 CNOT gates (instead of 15 CNOTs). As a result, the simulated NMR spectrum of the reduced Hamiltonian is significantly improved, as can be seen by comparing Fig. 4 with Fig. 6. For completeness, we have also presented the simulation with depolarising noise, and we see that the resulting spectrum agrees well with the simulated one on IBMQ cloud.
Deriving the effective line width
Having seen that the proposed effective linewidth Eq. 13 agrees well with the simulated spectrums on IBM’s quantum computers, we now outline how to obtain such a linewidth $\Gamma_{\rm eff}$. This derivation is based on the assumption that the error probability $\epsilon_{kl}$ is caused by incoherent errors. As shown above this seems to be a good approximation, but it is also possible to turn coherent error into incoherent error using randomized compiling without gate overhead[48] and hence the applicability of our assumption can be generalized.
Typically each single qubit is characterized by the parameters $T_{1}$ and $T_{2}^{*}$ [49, 50]. The former is the damping time scale, i.e., the time scale after which a qubit in an excited state will decay to its ground state.The second time scale $T_{2}^{*}$ is pure dephasing time, which refers to the time scale, after which the qubit’s phase is lost (when removing the contribution from damping). However, it has been shown in Ref. [51], that large angle gates drastically modify the form of such noise processes, leading to a form close to depolarizing noise [51]. And hence as all the native gates for IBM’s quantum processors are corresponding to large angle rotations (except the $R_{z}$ which is a virtual gate), in our modeling of IBM’s quantum computing devices, we model the gate errors as depolarizing noise. This noise model is also consistent with noise spectroscopy performed on the IBM devices [52]. However, even if noise would correspond to decay or pure dephasing this would not fundamentally change the effective line width. Incoherent errors are in general described by using a Lindblad superoperator $\mathcal{L}_{kl}$, with the noise strength characterized by decoherence rates $\gamma^{i,kl}_{\textrm{depol}}$ acting on qubit $i$. Since the error probabilities are given for each gate, it is best to model the effect of noise as a depolarising channel that acts after each gate. That is why we used the indices $k$ and $l$ for the Lindblad superoperator and the decoherence rate since each of these operators is associated with a unitary operations $\hat{U}_{kl}$.
To determine the decoherence rates of the depolarising noise after each gate $\hat{U}_{kl}$, we use the universal connection between incoherent noise and gate fidelity [53], which yields
$\epsilon_{kl}=\frac{3d}{4(d+1)}t_{\rm gate,kl}\sum_{i}^{n}\gamma^{i,kl}_{\textrm{depol}},$ (14)
where $n$ is the number of qubits the gate is acting on, $d=2^{n}$ is the dimension of the space and $t_{\rm gate,kl}$ is the gate execution time. The gate execution time is not necessarily exactly known, but as we will see it is also not necessarily needed. The relevant part of Eq. 14 is to show that there is a connection between error probabilities and decoherence rates. This is what we will use to derive $\Gamma_{\rm eff}$.
Since we think about a noisy system, we consider the time evolution of the density matrix $\rho$ and we can write the application of one Trotter step as
$\rho(\tau_{0}+\tau)=\prod_{l}e^{\mathcal{L}_{kl}t_{\rm gate,kl}}\mathbf{U}_{kl}\rho(\tau_{0})$ (15)
with $\mathbf{U}_{kl}\rho(\tau_{0})=\hat{U}_{kl}\rho(\tau_{0})\hat{U}_{kl}^{\dagger}$. It should be noted that we used the effective time of the Trotterized time evolution $\tau$ to indicate the time evolution of the density matrix. However, the effect of the noise is of course related to the physical gate execution time $t_{\rm gate,kl}$. For a Trotterized time evolution it is possible to efficiently commute all unitary operations in Eq. 15 to the left and all incoherent noise terms to the right. If we then rescale all depolarizing rates to
$\tilde{\gamma}^{i,kl}_{\textrm{depol}}\tau=\gamma^{i,kl}_{\textrm{depol}}t_{\rm gate,kl}$ (16)
we can write the master equation as
$\rho(\tau_{0}+\tau)=e^{([H,\bullet]+\mathcal{L}_{\rm incoherent})\tau}\rho(\tau_{0})$ (17)
where $e^{[H,\bullet]}$ represents the coherent time evolution we wanted to create using the Trotterization algorithm and $e^{\mathcal{L}_{\rm incoherent}\tau}$ contains all incoherent effects of the depolarizing noise. Details on the exact derivation of $\mathcal{L}_{\rm incoherent}$ can be found in Ref. [32]. Additionally, simple examples for the form of $\mathcal{L}_{\rm incoherent}$ can be found in Appendix B of Ref. [54]. However, the relevant transformations change the structure of the noise operators, but leaves the total sum of all incoherent processes $\sum_{ikl}\tilde{\gamma}^{i,kl}_{\textrm{depol}}$ unchanged. It should be noted that all depolarising rates $\tilde{\gamma}^{i,kl}_{\textrm{depol}}$ in $e^{\mathcal{L}_{\rm incoherent}\tau}$ are now connected to the error probabilities as
$\sum_{kl}\epsilon_{kl}\approx\frac{1}{2}\tau\sum_{kl}\sum_{i}^{n}\tilde{\gamma}^{i,kl}_{\textrm{depol}},$ (18)
where we have simplified the prefactor in Eq. 14 to $1/2$, which is a fair approximation if only one- and two-qubit gates are involved. Now if we assume that either a SWAP algorithm is used or that error probabilities are similar for all qubits used we get an effective averaged single spin decoherence rate by dividing the total sum of all rates by the number of spins, resulting in the full-width-half-maximum of
$\Gamma_{\rm eff}=\frac{\sum_{kl}\epsilon_{kl}}{N\tau},$ (19)
which effectively is the sum over all incoherent rates divided by the number of qubits. Such an effective decoherence rate can be also obtained from the equation of motion of correlation function within the effective open quantum system description, as is presented in Appendix E.
Discrete noise model in quantum circuits
The imperfections of the qubits typically results in incoherent errors, i.e., qubit damping [49], dephasing [50], or depolarization after the application of gates [32, 51]. Additionally, another source of error is the erroneous unitary operation implemented by the gates. Such coherent error can be interpreted as effective additional unitary gate. Moreover, leakage out of the computational basis of the qubit [55], and a fast return back to mentioned basis may be mapped to incoherent errors. In this section, we expand the noise model beyond incoherent errors. This allows us to reproduce most of the features of the shown spectrums simulated on IBM quantum processors.
Figure 7: (a) Comparison of the NMR spectrum of cis-3-chloroacrylic acid simulated on IBM Perth (solid orange curve) with on-gate noise model emulation (dashed blue curve) for 81 Trotter steps of size $\tau=0.01$. (b) Comparison of the NMR spectrum of 1,2,4-trichlorobenzene ($C_{6}H_{3}Cl_{3}$) simulated on IBM Perth, shown as orange solid curve, with both the numerical model emulation including only depolarising noise shown as the dashed blue curve, and a noise model including also coherent errors illustrated by the dotted magenta curve, for 21 Trotter steps of size $\tau=0.005$. In the latter noise model, we have assumed a coherent Z-error with angle $\phi_{z}=-0.027$, after each gate operation. After each two-qubit gate operation (here, CNOT operations), we have assumed a coherent error of the form $R_{x}$, with angle $\phi_{x}=0.05,-0.047$ on the control and target qubits, respectively, and an additional error $R_{z}(-0.05)$ solely on the control qubit.
Figure 7(a) illustrates the comparison of the simulation which includes only depolarising noise according to Eq. (14) with the IBM results for cis-3-chloroacrylic acid. As it is shown, the agreement is rather good. The peak positions show slight differences and also vary between different runs, see Appendix D, which is a manifestation of the effect of quasistatic coherent errors. Also the peak widths can vary between different runs.
Figure 7(b) shows the comparison of the simulation which includes only depolarising noise (blue dashed lines) with the IBM results for the case of 1,2,4-trichlorobenzene. This system involves more interaction terms and hence more gates in the circuit for time evolution [in total 15 CNOTs per Trotter step, in comparison to 3 CNOTs in case (a)]. The latter results in more prominant impact of coherent errors and hence a considerable deviation from the simulation with including only depolarising noise. As demonstrated, in the second numerical simulation (dotted lines), where we include coherent errors by applying an additional $R_{z}$ and $R_{x}$ on each involved gate (determined by a fitting routine), we can better reproduce the results obtained on IBM devices by such an inclusion of coherent errors. We use angles whose magnitudes are consistent with the noise spectroscopy of Ref. [52] and correspond to gate errors that are smaller than given in the calibration data. As illustrated, with the proper choice of angles corresponding to coherent errors, we are able to reproduce the main features of the simulated spectrum on IBM. This demonstrates that also coherent errors can effectively broaden the spectrum in large-depth circuits.
@@ Acknowledgements.
This work was supported by the German Federal Ministry of Education and Research, through projects QSolid (13N16155) and Q-Exa (13N16065). We would like to thank Peter Pinski, Sonia Alvarez, Julius Kleine Büning, and Jesse A. Vaitkus for insightful input and fruitful discussions.
@@ Appendix A Time evolution of Magnetization
In this section, we look at the time evolution of the correlation function simulated on IBM quantum computing platforms. Figure 8 depicts the correlation function $\text{Tr}\{S^{z}_{\rm tot}(t)S^{z}_{\rm tot}\}$ as a function of time, indicating the decay of correlation function with time. In the NMR experiment the spectrum is measured at sufficiently long times after the application of the magnetic field such that the system reaches a state in which average magnetization vanishes and hence the correlation function. As is shown in Fig. 8 there remain some small oscillations around zero at long times due to the limitation of number of Trotter steps. As the spectrum is obtained via the Fourier transform of the correlation function, see Eq. (10), the frequency interval is inversely proportional to the number of Trotter steps. Thereby we added additional zeros to the correlation functions in order to be able to obtain the spectrum on a finer frequency grid. Fig. 9 presents the comparison of the spectrum with and without adding such additional zeros for one of the simulations on IBM’s devices.
Figure 8: Time evolution of the correlation function of the cis-3-chloroacrylic acid, simulated on two different devices from IBM quantum computers, namely IBM Perth (orange curve) and IBM Nairobi (green curve), with 81 Trotter steps of size $\tau=0.01$.
@@ Appendix B The symmetrized NMR spectra
We discuss the symmetry breaking that occurs in the simulations as an artifact of noise of NISQ devices. In particular in Fig. 10, we present the magnetization in $y$ direction simulated on IonQ’s Aria. Note that due to the symmetry of the Hamiltonian of cis-3-chloroacrylic acid, see Eq. 2, we expect $\langle S^{0}_{y}\rangle=-\langle S^{1}_{y}\rangle$, such that the total magnetization in the $y$ direction vanishes. However, due to the presence of noise, such a symmetry is numerically broken. This is a manifestation of coherent errors [47] resulting in additional disordered effective Hamiltonian terms that break the mentioned symmetry. In particular, in additional numerical simulations we find that the spectroscopy on quantum computer is especially fragile against Hamiltonian disorder terms $\propto YZ$.
Figure 10: Time evolution of the magnetization in the $y$ direction for cis-3-chloroacrylic acid with Hamiltonian Eq. 2, simulated on IonQ Aria, with 61 Trotter steps of size $\tau=0.01$. The contribution from each spin to the magnetization (orange and blue curve) is compared with the total one (black curve).
Figure 11: Comparison of the symmetrized spectrum of the cis-3-chloroacrylic acid, simulated on IonQ Aria (dash-dotted blue curve) with 61 Trotter steps of size $\tau=0.01$, with exact diagonalization results (black solid curve).
@@ Appendix C The Trotter error
For a Trotterized time evolution, it is advantageous to increase the trotter step size to be able to simulate longer times, and also to decrease the impact of noise, as discussed in the Method section. However, by increasing the Trotter step size, we get an error from the Trotterization of the time evolution. Fig. 12 depicts the occurrence of such an error in the form of shifts of the resonances. Note that in the simulation with an on-gate depolarizing noise, we can reproduce such shifts to the actual resonances.
Figure 12: Simulation of NMR spectrum of cis-3-chloroacrylic acid on IBM Perth with 81 Trotter steps of size $\tau=0.02$ shown by orange curve, compared with exact results (black curve) as well as simulation with on-gate depolarizig noise model (dashed blue curve).
@@ Appendix D Additional results
In this section, we present further simulations of the NMR spectrum of cis-3-chloroacrylic acid on IBM’s processors. It is interesting to note that the deviations of the predicted resonances from the exact result varies in simulations performed on different days, which can be a manifestation of quasistatic coherent noise.
@@
Figure 13: Comparison of the simulation of NMR spectrum of cis-3-chloroacrylic acid on IBM Perth on three different days (shown with different colors) with exact results (black curve). For the simulation we performed a time evolution with 81 Trotter steps of size $\tau=0.01$.
Appendix E Effective decoherence rate
In this section, we outline the details of extracting the effective decoherence rate from the equation of motion of spin-spin correlation function. As has been discussed in the main text such an effective decoherence rate allows us to find the precision with which the NMR spectrum is simulated on quantum chips. To this end, we first aim to find the open quantum system description of applying the Trotterized time evolution on a noisy quantum chip. Such an open quantum system evolves according to master equation
$\dot{\rho}=\mathcal{L}\rho$ $=-i[H_{\textrm{gates}}(t),\rho]$
$+\sum_{\alpha}\gamma_{\alpha,\beta}\left(O_{\alpha}\rho O_{\beta}^{\dagger}-\frac{1}{2}\{O_{\alpha}O_{\beta}^{\dagger},\rho\}\right),$ (20)
where the matrix elements can be found from the noise mapping procedure explained in Ref. [32], see Fig. 14.
Figure 14: The sketch of noise mapping procedure outlined in Ref. [32], where the noisy qubits and noisy erroneous gates involved in the implementation of the Trotterized time evolution result in an effective open quantum system described by Eq. 20.
The spin operator like $\sigma^{\alpha}_{i}$, with $\alpha=x,y,z$ of the system then satisfies the following equation of motion
$\langle\dot{\sigma^{\alpha}_{i}}\rangle=$ $\text{Tr}\{O\dot{\rho}\}=-i\text{Tr}\left\{[\sigma^{\alpha}_{i},H_{\textrm{gates}}(t)]\rho\right\}$
$+\sum_{\alpha}\gamma_{\alpha}\text{Tr}\left\{\sigma^{\alpha}_{i}\left(O_{\alpha}\rho O_{\alpha}^{\dagger}-\frac{1}{2}\{O_{\alpha}O_{\alpha}^{\dagger},\rho\}\right)\right\}.$ (21)
Working in the basis of Pauli operators for noise operators, we can show that
$\langle\dot{\sigma^{\alpha}_{i}}\rangle=\sum_{j}M_{ij}\langle\sigma^{\alpha}_{i}\rangle.$ (22)
Moreover the quantum regression theorem implies that
$\langle S^{z/y}_{\rm tot}(t)S^{z}_{\rm tot}\rangle$ $=\text{Tr}\left\{e^{iHt}S^{z/y}_{\rm tot}e^{-iHt}S^{z}_{\rm tot}\rho(0)\times\rho_{\rm B}\right\}$
$=\text{Tr}_{\rm s}\left\{S^{z/y}_{\rm tot}\text{Tr}_{\rm B}\left\{e^{-iHt}S^{z}_{\rm tot}\rho(0)\times\rho_{\rm B}e^{iHt}\right\}\right\}$
$=\text{Tr}_{\rm s}\left\{S^{z/y}_{\rm tot}\mathcal{L}\left(S^{z}_{\rm tot}\rho(0)\right)\right\}$ (23)
where we have assumed the density operator of the total system, i.e., system togeather with bath, is $\rho(0)\times\rho_{\rm B}$. To better outline the implication of the regression theorem on the equation of motion of correlation functions, let us consider the case of one spin system $H=\epsilon\sigma_{x}$ with depopolarizing noise with rate $\Gamma_{\rm dep}$. We can show that using cyclic properties of trace, Eq.21 results in
$\langle\dot{\sigma^{z}}\rangle$ $=2\epsilon\langle\sigma^{y}\rangle-4\Gamma_{\rm dep}\langle\sigma^{z}\rangle,$
$\langle\dot{\sigma^{y}}\rangle$ $=-2\epsilon\langle\sigma^{z}\rangle-4\Gamma_{\rm dep}\langle\sigma^{y}\rangle,$
$\langle\dot{\sigma^{x}}\rangle$ $=2\epsilon\langle\sigma^{x}\rangle,$ (24)
and for the correlation functions, it follows
$\frac{d}{dt}\langle\sigma^{z}(t)\sigma^{z}(0)\rangle$ $=-2i\epsilon\langle i\sigma_{y}(t)\sigma_{z}(0)\rangle-4\Gamma_{\rm dep}\langle\sigma_{z}(t)\sigma_{z}(0)\rangle$ (25)
$\frac{d}{dt}\langle i\sigma^{y}(t)\sigma^{z}(0)\rangle$ $=-2i\epsilon\langle i\sigma_{z}(t)\sigma_{z}(0)\rangle-4\Gamma_{\rm dep}\langle i\sigma_{y}(t)\sigma_{z}(0)\rangle$ (26)
we can decouple the equations of motion for the linear combinations $\sigma^{z}\pm i\sigma^{y}$,
$\frac{d}{dt}\langle[\sigma^{z}(t)\pm i\sigma^{y}(t)]\sigma^{z}(0)\rangle$
$=\left(\mp 2i\epsilon-4\Gamma_{\rm dep}\right)\langle[\sigma^{z}(t)\pm i\sigma^{y}(t)]\sigma^{z}(0)\rangle$ (27)
which results in
$\langle\sigma^{z}(t)\sigma^{z}(0)\rangle=e^{-4\Gamma_{\rm dep}t}\bigg[$ $\cos(2\epsilon t)\langle\sigma^{z}(0)\sigma^{z}(0)\rangle$
$+\sin(2\epsilon)\langle\sigma^{y}(t)\sigma^{z}(0)\rangle\bigg].$ (28)
Figure 15: The NMR spectrum of $CHCl_{3}$ that has one spin for 1000 trotter steps of trotter time step $\tau=0.005$. The shaded area shows the broadening $4\Gamma_{\rm dep}$ that indicates the decay of the correlation functions, see Eq. 28.
For larger systems, for liquid phase hamiltonian of Eq. 1, and depolarizing rate of strength $\Gamma_{\rm dep}$, we can show that
$\langle\dot{\sigma}^{z}_{\rm tot}\rangle=2i\sum_{j}\epsilon_{j}\langle\sigma^{y}_{i}\rangle-4\Gamma_{\rm dep}\langle\sigma^{z}_{\rm tot}\rangle,$ (29)
$\langle\dot{\sigma}^{y}_{i}\rangle=$ $-2i\epsilon_{i}\langle\sigma^{z}_{i}\rangle-4\Gamma_{\rm dep}\langle\sigma^{y}_{i}\rangle$
$+2iJ\left[\langle\sigma^{x}_{i}\sigma^{z}_{j\neq i}\rangle-\langle\sigma^{x}_{j\neq i}\sigma_{z}^{i}\rangle\right],$ (30)
that indicates how the spin-spin coupling results in the coupling of equation of motion of spin operators with correlation functions. Ideally one has to solve the full set of coupled equations and find the decay of the correlation function which defines the effective decoherence rate. Note that while imaginary part of the right-hand side of Eqs. 29 and 30, that results in oscillatory behaviour are coupled, the real part, resulting in the decay seems quite simple in the case of depolarizing noise. As in a generic case, finding the full solution of Eqs. 29 and 30 is not feasible, we assume that such a coupling does not impact the decay rate, i.e., the effective decoherence rate. In the following we investigate to what extent such an assumption is valid by looking at the simulated spectrums of systems of various sizes.
Another complication that arises for larger systems, is that the effective open quantum system description is more complex than having a pure depolarizing noise. Note that even if we assume that the main processes describing the decoherence events for qubits are depolarizing noise, the presence of large angle gates, results in various complex noise terms. For instance, for the case of cis-3-chloroacrylic acid, the open quantum system consists of single and two particle noise operators as shown in figure 16.
Figure 16: The matrix representation of decoherence rates $\gamma_{\alpha,\beta}$, see Eq. 20, corresponding to a trotterized time evolution with $\tau=0.0025$ for cis-3-chloroacrylic acid.
For instance the noise term ”iXiX” results in the following contribution
$\gamma_{iXiX}\left(\text{Tr}\{\sigma^{z}_{k}\sigma^{x}_{i}\sigma^{x}_{i}\}-\frac{1}{2}\text{Tr}\left\{\sigma^{z}_{k}\{\sigma^{x}_{i}\sigma^{x}_{i},\rho\}\right\}\right)$
$=-2i\delta^{ik}\text{Tr}\{\sigma^{y}_{i}\sigma^{x}_{i}\}$ (31)
and similarly,
$\gamma_{iYiY}\left(\text{Tr}\left\{\sigma^{z}_{k}\sigma^{y}_{i}\sigma^{y}_{i}\right\}-\frac{1}{2}\text{Tr}\left\{\sigma^{z}_{k}\{\sigma^{y}_{i}\sigma^{y}_{i},\rho\}\right\}\right)$
$=-2i\delta^{ik}\text{Tr}\{\sigma^{x}_{i}\sigma^{y}_{i}\}$ (32)
therefore if $\gamma_{iXiX}=\gamma_{iYiY}$, as is the case , see Fig. 16, we have Eq. 31+Eq. 32 $4\gamma_{iXiX}\delta^{ik}\langle\sigma^{z}_{i}\rangle$.
Figure 17: The simulated NMR spectrum of cis-3-chloroacrylic acid that has 2 spins for 1000 trotter steps of various trotter time steps. The estimated broadening extracted from equation of motions is marked in different colors, while the definition Eq. 13 is the gray area.
Figure 18: The simulated NMR spectrum of $C_{6}H_{5}NO_{2}$ that has 5 spins for 200 trotter steps of trotter time step $\tau=0.001$. The estimated broadening extracted from equation of motions is marked with orang, while the definition Eq. 13 is shown with light blue.
For completeness we compare the estimated broadening extracted from equation of motion with the definition from the main text Eq. 13, where we consider a noisy simulation with on-gate depolarizing noise. As it is shown in Fig. 17 for a molecule with 2 spin, and in Fig. 18 with a molecule that has 5 spins, the comparison indicates that the estimated broadening in the main text is a good measure of the precision of the digital simulation of NMR spectrum on NISQ devices.